OpenStudy (anonymous):

integrate: 4dx/ (4-x^2)^(3/2)

OpenStudy (anonymous):

u= 4-x^2 du =-2xdx 4xdx = -2du integrate -2 (du/u^(3/2))

OpenStudy (anonymous):

@Algebraic! it is 4dx not 4xdx, your substitution is not correct.

OpenStudy (anonymous):

so it is...

OpenStudy (anonymous):

Correct. its 4 dx.

OpenStudy (anonymous):

@shawngomez36@gmail.com use trigonometric substitution method.

OpenStudy (turingtest):

for forms\[\sqrt{a^2-b^2x^2}\]let\[x=\frac ab\sin\theta\]

OpenStudy (anonymous):

i don't understand the u substitution method. can someone explain it?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\cos \theta = \frac{ \sqrt{4-x^2} }{ 2 }\] \[\cos ^{2} \theta = \frac{ 4-x^2 }{ 4 }\] \[\frac{ 1}{\cos ^{2} \theta} = \frac{ 4}{ 4-x^2 }\] \[\frac{ 1}{\cos ^{3} \theta} = \frac{ 4*2}{ (4-x^2)^{3/2} }\] \[\frac{ 1}{4*2\cos ^{3} \theta} = \frac{ 1}{ (4-x^2)^{3/2} }\] \[\frac{ x }{ 2 } = \sin \theta \] \[dx = 2 \cos \theta d \theta \] \[\frac{ 4dx }{ (4-x^2)^{3/2}} = \frac{ 4*2 \cos \theta d \theta }{ 4*2\cos ^{3} \theta }\] \[\frac{ d \theta }{ \cos ^{2} \theta } = \sec ^{2}\theta d\theta\]

OpenStudy (anonymous):

integrate to get tan(theta) which is x/(sqrt(4-x^2))